# PB-NBV: Efficient Projection-Based Next-Best-View Planning Framework for Reconstruction of Unknown Objects

### v1

@@ II Related Works
Finding an optimal 6-DOF viewpoint directly through optimization in space is very challenging, so researchers usually sample a number of candidate viewpoints and then choose the best one. Currently, common candidate viewpoint sampling strategies can generally be categorized into four types: random sampling[2, 3], predefine set[1, 6, 7], path-based sampling[13, 14, 15, 5], and scene-information-based sampling[8, 4]. The candidate viewpoint sampling strategy should be appropriately adjusted according to the current holistic data acquisition system. An appropriate strategy will enhance the convergence efficiency of the NBV planning framework.
@@ II-C2 Efficiency-focused strategies
The second category of strategies enhances the efficiency of ray-casting. Both [11] and [5] propose strategies to improve the efficiency of ray-casting itself. [22] uses GPU to accelerate ray-casting. Some deep learning strategies can avoid using ray-tracing algorithms and provide the best viewpoint at a fast speed. [6] and [9] use the network to predict the information gain for selecting the next perspective. [7] can infer all subsequent observation viewpoints by inputting only one frame of observation. However, generalizing such algorithms to other unknown objects that differ from their training set is challenging[8].
@@ III Methodology
III-A Framework Overview
The workflow of our projection-based NBV planning framework is shown in Fig 2. The robot arm first moves to a preset initial observation pose and then starts the iterative process of NBV planning.
@@ III-A Framework Overview
III-B Proposal of Candidate Viewpoints
@@
Considering the reachability of the robot arm, we set the sampling area of the candidate viewpoint to the partial hemisphere area above the turntable. The algorithm uses the base coordinates of the robot as the world coordinate system, as shown in Fig. 3. The radius of this hemisphere $R$ is the sum of the 3D camera’s working distance $d_{c}$ and half the diagonal length of the unknown object’s bounding box $d_{b}$. As the bounding box of the object changes during reconstruction, the size of the hemisphere is dynamically adjusted to accommodate the depth of field limitations of the 3D camera.
We selected $\alpha$ equidistant parallels from the candidate area and distributed $N$ candidate viewpoints proportionally along each parallel based on their lengths. All viewpoints are directed toward the center of the bounding box.
III-C Voxel Structure construction
@@
Each time the bounding box is updated, the newly added voxels are initially set to $V_{n}$, indicating voxels have not been scanned by the device. $V_{o}$ represents the surface structure of the unknown object, which is the voxel occupied by $P_{f}$. $V_{e}$ represents the free area in the bounding box, which is the set of voxels not occupied by $P_{f}$ after being scanned by the device. Once $V_{o}$ and $V_{e}$ are determined, they remain unchanged. $V_{u}$ represents the area where the surface structure of the unknown object may exist. These areas are not captured by the acquisition device because they are blocked by $V_{o}$. As shown in Fig. 4LABEL:sub@Voxel_struct_a, With each new addition of occupied voxels $V_{o}$, $V_{u}$ and $V_{e}$ will be generated accordingly.
We employ [24] to sample rays at equal angles within the FOV of the current viewpoint. In the voxels intersected by the ray, we distinguish between empty voxels $V_{e}$ and unknown voxels $V_{u}$ based on whether the current ray passes through $V_{o}$. The update of $V_{f}$ is carried out after the updates of $V_{u}$ and $V_{o}$. Any $v_{i}\in V_{u}$ is identified as a Frontier voxel if both empty and occupied voxels are present in its spatial neighborhood.
@@
The bounding box $\mathbf{B}$ is initialized to the bounding box of $V_{o}$ in the first frame. Then, $\mathbf{B}$ is expanded along the direction of the current viewpoint, doubling its diagonal length. In subsequent updates, a sphere $s_{i}$ with a radius of $\gamma$ is generated around each voxel $v_{i}\in V_{f}$, and the size of $\mathbf{B}$ is expanded to cover $V_{o}$, $V_{u}$ and $s_{i}$. The result is shown in Fig. 4LABEL:sub@Voxel_struct_b.
III-D Ellipsoid Representation
@@
Occupied voxels and frontier voxels will directly affect the viewpoint quality evaluation results. Therefore, after each update of $V_{o}$ and $V_{f}$, we will initialize $T$ 3D Gaussian distributions for each type of voxel ($V_{o}$ and $V_{f}$). We utilize the expectation-maximization (EM) algorithm to train a Gaussian distribution, which allows us to obtain the clustering results $C_{o}$ and $C_{f}$ after categorizing these two types of voxels.
@@
III-E Projection-based Viewpoint Quality Evaluation Function
@@
III-E1 Ellipsoid observation weight calculation
@@
III-E2 Ellipsoid weighted projection calculation
@@
[…]
@@
Where $Q^{*}$ is the dual matrix of $Q$, $\Phi^{*}$ is the dual matrix of $\Phi$. If any matrix $M$ is a reversible matrix, its dual matrix is $M^{*}=M^{-1}$. The camera matrix $P$ can be calculated by the internal parameter $K$ of the current 3D camera, and the camera’s pose $H$ in the world coordinate system[28].
@@
III-E3 Viewpoint quality evaluation function
@@
Figure 8: Simulation environments for the comparative experiments. (a) Gazebo simulation environment. (b) Objects to be tested, 55 in total.
III-F Global partitioning strategy
@@
Figure 9: The simulation results of each iteration in the comparison experiment for 55 objects. (a) Average coverage. (b) Average compute time.
@@ IV Experiments
IV-A Simulation experiment
To evaluate the performance of our framework, we constructed a simulation environment using the Gazebo platform on the Ubuntu operating system. Computations are conducted on a device with an i7-14700f processor. The simulation environment is configured with a single RealSense camera and includes 55 objects selected from the Stanford 3D Scanning Repository(graphics.stanford.edu/data/3Dscanrep/), Linemod, and HomebrewedDB(bop.felk.cvut.cz/datasets/). The simulation environment is depicted in Fig. 8.
The Simulation experiment is based on two evaluation metrics: (1) Point Cloud Coverage: Sampling 10,000 points from the model file, traversing all model points, and calculating the proportion of model points that have a point in the input point cloud within a distance of 0.005m. (2) Computational Time: Measuring the time taken by the NBV planner to provide the next best view after receiving a frame input.
@@ IV-A1 Comparison experiments
In Comparison experiments, our framework is compared with the voxel-based MCMF[2] (completeness-focused) NBV-net[6] and SCVP[7] (efficiency-focused), and the point cloud-based SEE[8].
In our experiments, both our framework and MCMF use an Octomap resolution of 0.03 meters. NBV-Net and SCVP, needing 32×32×32 voxels, use a resolution of 0.006 meters. Our framework, MCMF, SEE, and NBV-Net have 800 candidate viewpoints, while SCVP has 32. Both SCVP and NBV-Net adopt the pre-trained models provided by the SCVP open-source project. Our global partition number $\beta$ is 4, and $T_{max}$ is 10. Since there is no issue with the robotic arm being unreachable in the simulation, we set the candidate viewpoint sampling space to a complete sphere. In this experiment, our framework processed an average of 2123.40 voxels and an average of 11.06 ellipsoids (including occupied and frontier).
TABLE I: Comparison experiment result
Framework SEE SCVP NBVNET MCMF Ours
Coverage (%) 73.24 86.08 75.92 88.46 98.41
Time (s) 14.56 1.45 6.70 27.16 5.10
All frameworks receive data from the Realsense camera and adjust the camera’s pose based on their NBV outputs for iteration, starting with the same initial pose. Each framework performs ten iterations, with point cloud coverage and computational time analyzed as key metrics.
Fig.9 shows that our framework outperforms other frameworks in the 55 objects, not only with the highest point cloud coverage but also with lower computational time. Tab. I presents the final average coverage and the computational time per iteration for the 55 objects. Compared to MCMF, the second-highest coverage framework, our coverage is 1.12 times higher, while the computation time is only 0.18 times that of MCMF, which is lower than the neural network-based framework NBV-Net.
@@ IV-A2 Ablation experiment
We set $\beta$ to 1 (disabling the global partition strategy) and varied $T_{max}$ as 5, 10, 20, and 40, comparing them with the random selection strategy. The results shown in Fig.10LABEL:sub@Albation_T_max_CoverageLABEL:sub@Albation_T_max_Time indicate that the strategy is not sensitive to parameter selection, as different $T_{max}$ values have minimal impact on convergence efficiency, and increasing $T_{max}$ leads to a higher computational burden.
When $T_{max}$ is set to 10, we performed a comparison for global partition numbers $\beta$ of 1, 2, 4, and 8. The results show that a higher partition number improves the final convergence efficiency but at the cost of a slower convergence rate.
To verify the stability of our framework under different initial conditions, we tested it with $T_{max}=10$ and $\beta=4$ using four different initial viewpoints: [0.9,0,0], [0,0.9,0], [0,0,0.9], and [0,0.63,0.63]. The results, as shown in Fig.10LABEL:sub@Albation_Init_Pose_CoverageLABEL:sub@Albation_Init_Pose_Time, indicate that our framework is stable across different initial values.
Figure 10: Ablation experiment results for 55 objects in the simulation environment. (a)(c)(e) Average coverage per iteration. (b)(d)(f) Average compute time per iteration.
@@ IV-B Real-world experiments
As shown in Fig. 1, our experimental platform includes a self-made structured light 3D camera, a robotic arm with six degrees of freedom and a working range of 800 mm. and a 360-degree rotatable turntable to cover the entire hemispherical space of the object. The NBV planning framework and control programs run on a laptop with an i9-14700HX processor. The experiment used an Octomap resolution of 0.01m, with the number of partitions $\beta$ set to 4, the maximum number of fitted ellipsoids $T_{m}ax$ set to 50, the initial observation pose as shown in Fig. 1, and the number of iterations is 10. Reconstruction experiments on four objects of different sizes showed that our framework could achieve good results when reconstructing different objects, proving its feasibility.
Figure 11: The models used in the real-world experiment and their reconstruction results, from top to bottom: statue, vase, rhino and dog.
TABLE II: Object point cloud size
Object name statue vase rhino dog
point cloud size 9655 3013 3857 2684
Since there is no ground truth for objects in real-world experiments, we adopted the SEE[8] by treating the final point cloud from the iterations as the ground truth to assess the convergence efficiency of the NBV process. The iteration process is shown in Fig. 12.
Tab. II presents the number of points in the registered point cloud after filtering with voxels of 0.005m edge length, which is used to indicate the size of the object.
@@
Fig. 12 and Tab. II show that objects like the vase and dog have lower computation times. Complex objects like the statue and rhino need more ellipsoids, increasing computation times. Although the point cloud size of the statue is more than twice that of the rhino, the computation time is similar, indicating that the computational cost of the framework is not significantly related to the object volume size.

### v2

@@ II-B Candidate Viewpoint Proposal
Finding an optimal 6-DOF viewpoint directly through optimization in space is very challenging, so researchers usually sample a number of candidate viewpoints and then choose the best one. Currently, common candidate viewpoint sampling strategies can generally be categorized into four types: random sampling[2, 3], predefined set[1, 6, 7], path-based sampling[13, 14, 15, 5], and scene-information-based sampling[8, 4]. The candidate viewpoint sampling strategy should be appropriately adjusted according to the current holistic data acquisition system. An appropriate strategy will enhance the convergence efficiency of the NBV planning framework.
@@ II-C2 Efficiency-focused strategies
The second category of strategies enhances the efficiency of ray-casting. Both [11] and [5] propose strategies to improve the efficiency of ray-casting itself. [22] uses the GPU to accelerate ray-casting. Some deep learning strategies can avoid using ray-tracing algorithms and provide the best viewpoint at a fast speed. [6] and [9] use the network to predict the information gain for selecting the next perspective. [7] can infer all subsequent observation viewpoints by inputting only one frame of observation. However, generalizing such algorithms to other unknown objects that differ from their training set is challenging [8].
@@ III Methodology
The workflow of our projection-based NBV planning framework is shown in Fig. 2. The robot arm first moves to a preset initial observation pose and then starts the iterative process of NBV planning.
@@
III-A Proposal of Candidate Viewpoints
@@
Considering the reachability of the robot arm, we set the sampling area of the candidate viewpoint to the partial hemisphere area above the turntable. The algorithm uses the base coordinates of the robot as the world coordinate system, as shown in Fig. 3. The radius of this hemisphere $R$ is the sum of the 3D camera’s working distance $d_{c}$ and half the diagonal length of the unknown object’s bounding box $d_{b}$. Inspired by unknown-target scanning strategies from [24, 25], the hemisphere dynamically resizes with the object’s bounding box to fit the 3D camera’s depth of field. We selected $\alpha$ equidistant parallels from the candidate area and distributed $N$ candidate viewpoints proportionally along each parallel based on their lengths. All viewpoints are directed toward the center of the bounding box.
III-B Voxel Structure Construction
@@
All newly generated voxels are initialized to state $V_{n}$. Upon completion of sensor observation, multiple ray bundles are cast toward the target region through ray-travel[26], with voxel states updated according to the following rules: 1) If point cloud data $P_{f}$ resides within a $V_{n}$ voxel, the voxel is updated to $V_{o}$, indicating an occupied region of the target surface; 2) When a $V_{o}$ voxel exists along the ray path, all $V_{n}$ voxels penetrated by the ray behind this $V_{o}$ voxel are labeled as $V_{u}$, representing occlusion-induced uncertainty; 3) All $V_{n}$ voxels along the ray path prior to encountering any $V_{o}$ voxel are updated to $V_{e}$, denoting observed free space. 4) The update of $V_{f}$ is carried out after the updates of $V_{u}$ and $V_{o}$. Any $v_{i}\in V_{u}$ is identified as a Frontier voxel if both empty and occupied voxels are present in its spatial neighborhood.
@@
The bounding box $\mathbf{B}$ is initialized to the bounding box of $V_{o}$ in the first frame. Then, $\mathbf{B}$ is expanded along the direction of the current viewpoint, doubling its diagonal length. In subsequent updates, a sphere $s_{i}$ with a radius of $\gamma$ is generated around each voxel $v_{i}\in V_{f}$, and the size of $\mathbf{B}$ is expanded to cover $V_{o}$, $V_{u}$, and $s_{i}$. The result is shown in Fig. 4LABEL:sub@Voxel_struct_b.
III-C Ellipsoid Representation
@@
Occupied voxels and frontier voxels will directly affect the viewpoint quality evaluation results. Therefore, after each update of $V_{o}$ and $V_{f}$, we will initialize $T$ with 3D Gaussian distributions for each type of voxel ($V_{o}$ and $V_{f}$). We utilize the expectation-maximization (EM) algorithm to train a Gaussian distribution, which allows us to obtain the clustering results $C_{o}$ and $C_{f}$ after categorizing these two types of voxels.
@@
III-D Projection-based Viewpoint Quality Evaluation Function
@@
III-D1 Ellipsoid observation weight calculation
@@
III-D2 Ellipsoid weighted projection calculation
@@
Where $X=[x\ y\ z\ 1]^{T}$. For any given ellipsoid, the ellipse it projects onto under the action of the camera projection matrix $P$ can be represented by the matrix description $\Phi$.
@@
For invertible quadric matrices $Q$ and $\Phi$, their dual representations satisfy $Q^{*}=Q^{-1}$ and $\Phi^{*}=\Phi^{-1}$, where the dual space encodes tangent plane (line) constraints. The camera matrix $P$ can be calculated by the internal parameter $K$ of the current 3D camera, and the camera’s pose $H$ in the world coordinate system[30].
@@
III-D3 Viewpoint quality evaluation function
@@
(a) Simulation environment
(b) 55 objects to be tested
Figure 8: Simulation environments in Gazebo.
III-E Global Partitioning Strategy
@@
(a) Coverage for each framework
(b) Compute time for each framework
Figure 9: Comparison experiment tests 55 objects in the simulation environment, with results showing the average performance of all objects per iteration.
@@ IV Experiments
IV-A Simulation Experiments
To evaluate the performance of our framework, we constructed a simulation environment using the Gazebo platform on the Ubuntu operating system. Computations are conducted on a CPU-only device with an i7-14700F processor. The simulation environment is configured with a single RealSense camera and includes 55 objects selected from the Stanford 3D Scanning Repository [^2] Stanford: https://graphics.stanford.edu/data/3Dscanrep/, Linemod, and HomebrewedDB [^3] Linemod and HomebrewedDB: https://bop.felk.cvut.cz/datasets/. The simulation environment is depicted in Fig. 8.
The Simulation experiment is based on two evaluation metrics: (1) Point Cloud Coverage: Sampling 10,000 points from the model file, traversing all model points, and calculating the proportion of model points that have a point in the input point cloud within a distance of 0.005m. (2) Computational Time: Measuring the computation time of the NBV planner per iteration, covering scene representation update and optimal viewpoint selection.
@@ IV-A1 Comparison experiments
In Comparison experiments, our framework (with full-sphere and hemisphere candidate viewpoints sampling area) will be compared with the voxel-based MCMF[2] (completeness-focused) NBV-Net[6] and SCVP[7] (efficiency-focused), and the point cloud-based SEE[8]. Key experimental parameters are specified in Tab. I, with remaining parameters set to their default values.
TABLE I: Performance Comparison
Frame Cov. (%) T. (s) Cand. Views Area Res. (m)
SEE [^4] The direction SEE’s viewpoint is defined by its planner, while the rest point to the object center. For resolution, SEE uses a point cloud density radius, while others use voxel size. 73.24 14.49 800 Hemi 0.03
SCVP [^5] Both SCVP and NBV-Net utilize the pre-trained model from the SCVP open-source project. As the model requires input voxels of size 32×32×32, we uniformly set the voxel resolution to 0.006 m and employed a fixed set of 32 candidate viewpoints 86.08 1.45 32 Hemi 0.006
NBVNET 85.40 6.91 32 Hemi 0.006
Ours-H 88.66 5.15 800 Hemi 0.03
MCMF 88.46 27.16 800 Full 0.03
Ours-F 98.41 5.10 800 Full 0.03
Cov.: Coverage; T.: Computation Time; Cand. Views: Candidate Views;
Res.: Resolution; Hemi: Hemisphere; Full: Full Sphere.
In this experiment, our global partition number $\beta$ is 4, and $T_{max}$ is 10, and our framework processed an average of 2123.40 voxels and an average of 11.06 ellipsoids (including occupied and frontier). All frameworks receive data from the RealSense camera and adjust the camera’s pose based on their NBV outputs for iteration, starting with the same initial pose. Each framework performs 10 iterations [^6] If any framework produces fewer than 10 views for an object, we duplicate the last view until we reach 10 views., with point cloud coverage and computational time analyzed as key metrics. This experiment uses 10 iterations as the benchmark, emphasizing the comparison of how efficiently different frameworks rapidly improve object scan coverage, rather than their ultimate convergence performance.
As illustrated in Fig. 9, comparative results show that our framework not only achieves the highest coverage but also maintains lower compute time compared to other benchmark frameworks. Table I presents the comparisons in different sampling regions, revealing that our framework consistently exhibits superior coverage and better computational efficiency in both hemispherical and full spherical sampling spaces. The experimental results demonstrate that the ellipsoid representation structure can efficiently and flexibly characterize the spatial distribution features of frontier regions. Integrated with a projection-based optimal viewpoint selection strategy, our framework enables rapid and accurate determination of the next best observation viewpoint, thereby significantly improving the scanning efficiency for unknown objects.
@@ IV-A2 Ablation experiment
We set $\beta$ to 1 (disabling the global partition strategy) and varied $T_{max}$ to 5, 10, 20, and 40, comparing them with the random selection strategy. The results shown in Fig.10LABEL:sub@Albation_T_max_CoverageLABEL:sub@Albation_T_max_Time indicate that the strategy is not sensitive to parameter selection, as different $T_{max}$ values have minimal impact on convergence efficiency, and increasing $T_{max}$ leads to a higher computational burden. When $T_{max}$ is set to 10, we performed a comparison for global partition numbers $\beta$ of 1, 2, 4, and 8. The results in Fig. 10LABEL:sub@Albation_Beta_CoverageLABEL:sub@Albation_Beta_Time show that a higher partition number improves the final convergence efficiency but at the cost of a slower convergence rate. To verify the stability of our framework under different initial conditions, we tested it with $T_{max}=10$ and $\beta=4$ using four different initial viewpoints: [0.9,0,0], [0,0.9,0], [0,0,0.9], and [0,0.63,0.63]. The results, as shown in Fig.10LABEL:sub@Albation_Init_Pose_CoverageLABEL:sub@Albation_Init_Pose_Time, indicate that our framework is stable across different initial values.
(a) Effect of $T_{max}$ on coverage
(b) Effect of $\beta$ on coverage
(c) Effect of initial pose on coverage
(d) Effect of $T_{max}$ on compute time
(e) Effect of $\beta$ on compute time
(f) Effect of initial pose on compute time
Figure 10: Ablation experiment tests 55 objects in the simulation environment, with results showing the average performance of all objects per iteration.
@@ IV-B Real-world Experiments
Figure 12: The models used in the real-world experiment and their reconstruction results, from top to bottom: statue (9,655 points), vase (3,013 points), rhino (3,857 points), and dog (2,684 points).
As shown in Fig. 1, our experimental platform includes a self-made structured light 3D camera, a six-axis robotic arm with 800 mm working envelope, and a 360-degree rotatable turntable to cover the entire hemispherical space of the object. The NBV planning framework and control programs run on a laptop with an i9-14700HX processor. The experiment used an Octomap resolution of 0.01m, with the number of partitions $\beta$ set to 4, the maximum number of fitted ellipsoids $T_{max}$ set to 50, the initial observation pose as shown in Fig. 1, and the number of iterations is 10. Reconstruction experiments on four objects of different sizes showed that our framework could achieve good results when reconstructing different objects, proving its feasibility. Fig. 12 presents the real-world experimental results, where the number of registered point clouds after voxel filtering (0.005m) is used to represent the object size.
Since there is no ground truth for objects in real-world experiments, we adopted the method from SEE[8] by treating the final point cloud from the iterations as the ground truth to assess the convergence efficiency of the NBV process. The iteration process is shown in Fig. 11.
Fig. 11 shows that objects like the vase and dog have lower computation times. Complex objects like the statue and rhino need more ellipsoids, increasing computation times. Although the point cloud size of the statue is more than twice that of the rhino, the computation time is similar, indicating that the computational cost of the framework is not significantly related to the object volume size.